# Rigidity of anti-de Sitter (2+1)-spacetimes with convex boundary near the Fuchsian locus

### v1

@@ 1.1. Motivation
A deep geometric result states that the shape of a convex body in Euclidean 3-space is completely determined by the intrinsic geometry of its boundary [48]. Using this fact as a foundation for our study, we aim to investigate, to which extent it can be generalized. Standard proofs of this result use significantly that a convex body has the topology of a 3-ball. However, it was observed that this phenomenon holds also in nontrivial topology, provided that we stay in the setting of homogeneous geometry. In the case of hyperbolic geometry, which is the richest of Riemannian homogeneous geometries in dimension 3, manifestations of this were established, e.g., in papers [54, 50].
In the present paper, we are concerned with anti-de Sitter geometry, which is a Lorentzian cousin of hyperbolic geometry. Namely, we deal with globally hyperbolic Cauchy compact anti-de Sitter (2+1)-spacetimes. They appear, e.g., in models of quantum gravity, see the fundamental paper of Witten [64]. They are reminiscent in their properties to quasi-Fuchsian hyperbolic 3-manifolds. In particular, it is conjectured that globally hyperbolic compact anti-de Sitter (2+1)-spacetimes with spacelike convex boundary are determined by the intrinsic geometry of the boundary. Every such spacetime has two boundary components and can be canonically extended beyond the boundary to a maximal spacetime. Fundamentally, the conjecture means that the geometry of the whole spacetime is determined by the intrinsic geometry of two of its spacelike slices provided that the slices are convex in the opposite directions. In this article, we prove the conjecture among spacetimes whose holonomy is close to Fuchsian and whose boundary is either smooth or polyhedral.
@@ 1.2. Parallels between hyperbolic and anti-de Sitter geometry
In the hyperbolic case, Thurston conjectured that the respective marked moduli space can be parameterized by the induced metrics on the boundary of the convex core as well as parameterized by the pleating laminations. These questions have two aspects: the realization part – to describe the image of the respective map from the marked moduli space, and the rigidity part – to show that the map is injective. For the first conjecture, the realization part follows, e.g., from the work of Labourie [40]; the rigidity is only known to hold for an open dense set, as shown in the work [50] of Prosanov. For the second conjecture, the image was described in [10] by Bonahon–Otal; the rigidity has been obtained only recently in [21] by Dular–Schlenker. An important earlier advance was made by Bonahon in [9], where he proved the rigidity in a neighborhood of the Fuchsian locus. Prior to that, it was only known in a related, though slightly different setting when $S$ is the once-punctured torus, due to Series [56].
@@ 1.3. Manifolds with boundary
As we saw, geometric rigidity questions are interconnected with the corresponding realization questions. In the case of Euclidean convex bodies, the relevant realization questions were answered by Nirenberg [46] in the smooth case, and by Alexandrov [1] in the polyhedral and, furthermore, in the general case.
@@
The existence part of this statement was proved by Tamburelli [59]. However, the uniqueness remains elusive. In fact, a key open question is, given such a spacetime $M$, whether any first-order deformation of the AdS structure on $M$ that leaves invariant the induced metric on the boundary is trivial. We will discuss this further below.
There is a polyhedral counterpart to this story. There are, however, some special peculiarities, since the most natural notion of polyhedral boundary (locally modeled on polyhedral sets) turns out to be insufficient to capture all the picture of what is happening. Some singularities, similar to those of pleating laminations, may appear, which obstruct the analysis of this case. We refer to paper [50] for a description of these phenomena in the hyperbolic setting. In that paper, the corresponding realization statement is obtained, and the rigidity is shown in some generic case. Luckily, these difficulties lay outside of the scope of the present article, since they happen away from the Fuchsian locus.
@@ 1.4. Our results
There exists a neighborhood $U$ of $\mathscr{M}_{0}$ in $\mathscr{M}$ and an open subset $U^{\prime}\subset\mathscr{S}\times\mathscr{S}$ such that $\mathscr{I}|_{U}$ is a homeomorphism from $U$ to $U^{\prime}$.
@@
For each of those AdS structures in $\mathscr{M}(V)$, the induced metric on $\partial M$ is hyperbolic, with cone-singularities of angle larger than $2\pi$ at the points of $V$. We consider $V_{\pm}$ as subsets of $S$ and denote by $\mathscr{S}(V_{\pm})$ the spaces of hyperbolic metrics with cone-singularities at $V_{\pm}$ respectively of angles $>2\pi$. We have the induced metric map $\mathscr{I}_{V}:\mathscr{M}(V)\rightarrow\mathscr{S}(V_{+})\times\mathscr{S}(V_{-})$. We then have the following results.
@@
There exists a neighborhood $U$ of $\mathscr{M}_{0}(V)$ in $\mathscr{M}(V)$ and an open subset $U^{\prime}\subset\mathscr{S}(V_{+})\times\mathscr{S}(V_{-})$ such that $\mathscr{I}_{V}|_{U}$ is a homeomorphism from $U$ to $U^{\prime}$.
@@ 1.5. A word on the proofs
Let $m\in\mathscr{M}$, and $(s_{+},s_{-})=\mathscr{I}(m)$ be the pair of the induced metrics on its boundary. Let $\dot{m}$ be a first-order deformation of $m$. Is the corresponding variation $(\dot{m}_{+},\dot{m}_{-})$ of the induced metrics on the boundary necessarily nonzero?
@@ 1.6. Related results
There is a vast body of related results in this area of geometry. Here we survey some of them that we have not mentioned yet, though our exposition does not aim to be exhaustive. For another version of such a survey we refer to [29] by Fillastre–Smith.
@@
Question 1 is also equivalent to a question of simultaneous isometric embeddings of two metrics on $S$ into GHMC AdS (2+1)-spacetimes. Another interesting problem is to attempt to embed only one metric into a spacetime with prescribed left or right metric in the sense of Mess. This problem in the smooth case was recently fully resolved in [18] by Chen–Schlenker. In the Minkowski setting, instead of left or right metric it is natural to prescribe the linear part of the holonomy. In the smooth case, this follows from the work [61] by Trapani–Valli, while in the polyhedral case, it is proved in [25] by Fillastre–Prosanov. Actually, we will employ these results in our proof. We note that analogous questions in the de Sitter and in the hyperbolic settings (when the left or right metric is replaced by a conformal structure at infinity) are fully open.
Acknowledgments. The first author is very grateful to François Fillastre, Thomas Körber, Andrea Seppi and Graham Smith for useful discussions and to Michael Eichmair for helpful comments and his support.
@@
$M$ 3-manifold diffeomorphic to $S\times\mathbb{R}$
$\mathcal{S}$ the space of Riemannian metrics on $S$ of curvature $<-1$
$\mathcal{M}$ the space of GHC AdS structures on $M$
@@
$\mathcal{R}$ $\pi_{1}S\rightarrow G$ up to conjugation
$\mathcal{E}_{\pm}$ equivariant strictly future-/past-convex embeddings of $S$ into $\mathbb{A}d\mathbb{S}^{3}$ up to isometry
$\mathcal{E}$ pairs of embeddings in $\mathcal{E}_{+}\times\mathcal{E}_{-}$ equivariant with respect to the same representation
$\mathcal{GC}_{\pm}(s)$ the spaces of Gauss–Codazzi fields for a metric $s$ on $S$
@@
We add the subscript $F$ for subspaces of Fuchsian objects, i.e., related to the representations of $\pi_{1}S$ into $G_{F}:={\rm PSL}(2,\mathbb{R})$. We add the superscript 00 when anti-de Sitter space is replaced by Minkowski space. We use the font $\mathscr{S}$, $\mathscr{M}$, etc., for the quotients of $\mathcal{S}$, $\mathcal{M}$, etc., by isotopies. We use the hat: $\widehat{\mathcal{R}}$, $\widehat{\mathcal{E}}$, etc., for spaces not up to conjugation/isometries.
@@ 2.1. Anti-de Sitter geometry
The induced metric is a Lorentzian metric of constant curvature $-1$.
We will, however, consider as the main model for anti-de Sitter 3-space the projective quotient of $\mathbb{H}^{2,1}$. We will denote it by $\mathbb{A}d\mathbb{S}^{3}$ and consider it as a subset of $\mathbb{RP}^{3}$. Recall that it is a minimal model space for anti-de Sitter geometry [15].
We also briefly recall the $\rm PSL(2,\mathbb{R})$-model for $\mathbb{A}d\mathbb{S}^{3}$, which is a special feature of dimension 3. Note that on the space of real 2$\times$2-matrices, the determinant is a quadratic form of signature $(2,2)$. Considering the polarization of the minus determinant as a bilinear form, we can identify this space with $\mathbb{R}^{2,2}$. Under this identification, $\rm SL(2,\mathbb{R})$ gets identified with $\mathbb{H}^{2,1}$, and $\rm PSL(2,\mathbb{R})$ gets identified with $\mathbb{A}d\mathbb{S}^{3}$.
This model admits a straightforward description of the isometry group of $\mathbb{A}d\mathbb{S}^{3}$. We are interested only in its identity component, which we denote by $G$. Since for $A,B,X\in\rm SL(2,\mathbb{R})$ we have $\det(X)=\det(AXB^{-1})$, the left and right multiplications by elements of $\rm PSL(2,\mathbb{R})$ are isometries. This produces a map $\rm PSL(2,\mathbb{R})\times\rm PSL(2,\mathbb{R})\rightarrow G$. A dimension count shows that this is actually an isomorphism
@@ 2.2. Quasi-Fuchsian representations
To proceed further, we need to recall the basics of Teichmüller theory. The Teichmüller space $\mathscr{T}$ is the space of isotopy classes of hyperbolic metrics on $S$. Denote by $\mathbb{H}^{2}$ the hyperbolic plane and by $G_{F}$ the identity component of its isometry group. Every hyperbolic metric on $S$ is a $(G_{F},\mathbb{H}^{2})$-structure in the sense of Thurston. For an introduction to Thurston’s approach to geometric structures (later we will also apply this framework to anti-de Sitter structures), we refer to Thurston’s notes [60] and an accompanying paper [16]. Provided that $S$ and $\mathbb{H}^{2}$ are considered oriented, the holonomies of orientation-preserving hyperbolic structures form a connected component in the representation space of $\pi_{1}S$ to $G_{F}$. We call such representations Fuchsian and denote by $\mathcal{R}_{F}$ the quotient of this component by conjugation of $G_{F}$, the component itself is denoted by ^$\widehat{\mathcal{R}}_{F}$. In terms of notation, we will not distinguish between representations and their classes. It is well-known that the map sending a hyperbolic structure to its holonomy provides an identification $\mathscr{T}\cong\mathcal{R}_{F}$. See, e.g., the book of Farb–Margalit [22] for an introduction.
Consider $G_{F}\cong\rm PSL(2,\mathbb{R})$ as the diagonal subgroup of $G\cong\rm PSL(2,\mathbb{R})\times\rm PSL(2,\mathbb{R})$. We call a representation $\rho:\pi_{1}S\rightarrow G$ quasi-Fuchsian if it is a deformation of a Fuchsian representation. By definition, quasi-Fuchsian representations form a connected component of a representation space. We denote by $\mathcal{R}$ the quotient by conjugation of $G$, the component itself is denoted by $\widehat{\mathcal{R}}$. The projections of a quasi-Fuchsian representation to both factors in (2.1) are discrete and faithful. Hence, we have ^^$\widehat{\mathcal{R}}\cong\widehat{\mathcal{R}}_{F}\times\widehat{\mathcal{R}}_{F}$ and
@@ 2.3. Globally hyperbolic anti-de Sitter spacetimes
An anti-de Sitter (2+1)-spacetime is a 3-manifold equipped with an anti-de Sitter metric, i.e., with a Lorentzian metric of curvature $-1$. Provided that it is oriented and time-oriented, it is equipped with a $(G,\mathbb{A}d\mathbb{S}^{3})$-structure in the sense of Thurston. Hence, every anti-de Sitter (2+1)-spacetime $M$ is equipped with a holonomy representation $\rho:\pi_{1}M\rightarrow G$ and a $\rho$-equivariant developing map $f:\widetilde{M}\rightarrow G$, which are determined up to an isometry of $\mathbb{A}d\mathbb{S}^{3}$. A (2+1)-spacetime is globally hyperbolic if it admits a Cauchy surface, i.e., a surface that is intersected exactly once by every inextensible causal curve. All Cauchy surfaces in a spacetime are homeomorphic to each other. See, e.g., the book [47] by O’Neill for an introduction to Lorentzian geometry. By a result of Geroch [31], any globally hyperbolic (2+1)-spacetime admits a parametrization $S\times\mathbb{R}$, in which all fibers $S\times\{r\}$ are Cauchy surfaces.
A globally hyperbolic spacetime $N$ is maximal if every isometric embedding $N\rightarrow N^{\prime}$ that sends a Cauchy surface of $N$ to a Cauchy surface of $N^{\prime}$ is onto. It is Cauchy compact if its Cauchy surfaces are compact. Recall that we refer to globally hyperbolic maximal Cauchy compact anti-de Sitter spacetimes as GHMC AdS for short. Any GHMC (2+1)-spacetime $N$ is diffeomorphic to $S\times\mathbb{R}$, and hence $\pi_{1}N=\pi_{1}S$. Let now $N$ be a manifold diffeomorphic to $S\times\mathbb{R}$, and let $\mathscr{N}$ be the space of isotopy classes of GHMC AdS structures on $N$, with the topology defined similarly to above. Mess showed in [44] that the holonomy of every GHMC AdS (2+1)-spacetime is quasi-Fuchsian. Furthermore, it follows from his work that the map $\mathscr{N}\rightarrow\mathcal{R}$ sending a GHMC AdS structure to its holonomy is a homeomorphism. Any spacelike surface $\Sigma\subset\mathbb{A}d\mathbb{S}^{3}$ invariant with respect to $\rho\in\mathcal{R}$ belongs to the domain of dependence $D(\rho)\subset\mathbb{A}d\mathbb{S}^{3}$, a convex open $\rho$-invariant set, whose quotient by $\rho$ produces the respective GHMC AdS spacetime.

### v2

@@ 1.1. Motivation
A deep geometric result states that the shape of a convex body in Euclidean 3-space is completely determined by the intrinsic geometry of its boundary [50]. Using this fact as a foundation for our study, we aim to investigate, to which extent it can be generalized. Standard proofs of this result rely significantly on the fact that a convex body has the topology of a 3-ball. However, it was observed that this phenomenon holds also in nontrivial topology, provided that we stay in the setting of homogeneous geometry. In the case of hyperbolic geometry, which is the richest of Riemannian homogeneous geometries in dimension 3, manifestations of this were established, e.g., in the papers [57, 51].
In the present paper, we are concerned with anti-de Sitter geometry, which is a Lorentzian cousin of hyperbolic geometry. Namely, we deal with globally hyperbolic Cauchy compact anti-de Sitter (2+1)-spacetimes. They appear, e.g., in models of quantum gravity, see the fundamental paper of Witten [68]. They are reminiscent in their properties to quasi-Fuchsian hyperbolic 3-manifolds. In particular, it is conjectured that globally hyperbolic compact anti-de Sitter (2+1)-spacetimes with spacelike convex boundary are determined by the intrinsic geometry of the boundary. See, e.g., [7, Questions 3.3–3.6]. Every such spacetime has two boundary components and can be canonically extended beyond the boundary to a maximal spacetime. Fundamentally, the conjecture means that the geometry of the whole spacetime is determined by the intrinsic geometry of two of its spacelike slices provided that the slices are convex in the opposite directions. In this article, we prove the conjecture among spacetimes whose holonomy is close to Fuchsian and whose boundary is either smooth or polyhedral. In the works [14, 54], some other partial cases of the conjecture are established, see below for details.
@@ 1.2. Parallels between hyperbolic and anti-de Sitter geometry
In the hyperbolic case, Thurston conjectured that the corresponding marked moduli space can be parameterized both by the induced metrics on the boundary of the convex core and by the pleating laminations. These questions have two aspects: the realization part – to describe the image of the respective map from the marked moduli space, and the rigidity part – to show that the map is injective. For the first conjecture, the realization part follows, e.g., from the work of Labourie [42]; the rigidity is only known to hold for an open dense set, as shown in the work [51] of Prosanov. For the second conjecture, the image was described in [11] by Bonahon–Otal; the rigidity has been obtained only recently in [23] by Dular–Schlenker. (An important earlier advance was made by Bonahon in [10], where he proved the rigidity in a neighborhood of the Fuchsian locus.) Prior to that, it was only known in a related, though slightly different setting when $S$ is the once-punctured torus, due to Series [60].
@@ 1.3. Manifolds with boundary
As we saw, geometric rigidity questions are interconnected with the corresponding realization questions. In the case of Euclidean convex bodies, the relevant realization questions were answered by Nirenberg [48] in the smooth case, and by Alexandrov [1] in the polyhedral and, furthermore, in the general case of metric structures of curvature $K\geq 0$ in the sense of Alexandrov, without any additional regularity assumption.
@@
The existence part of this statement was proved by Tamburelli [63]. However, the uniqueness remains elusive in general. In [14], Bonsante–Mondello–Schlenker showed the uniqueness for the case when $_{+}, ss_{-}$ have constant curvatures $\kappa_{+},\kappa_{-}<-1$ respectively, satisfying $\kappa_{+}=-\kappa_{-}/(\kappa_{-}+1)$.
In fact, a key open question is, given such a spacetime $M$, whether any first-order deformation of the AdS structure on $M$ that leaves invariant the induced metric on the boundary is trivial. We will discuss this further below.
There is a polyhedral counterpart to this story, when $s_{+},s_{-}$ are so-called cone-metrics. There are, however, some special peculiarities, since the most natural notion of polyhedral boundary (locally modeled on polyhedral sets) turns out to be insufficient to capture all the picture of what is happening. Some singularities, similar to those of pleating laminations, may appear, which obstruct the analysis of this case. We refer to [54] of Prosanov for a description, and to [51] also of Prosanov for a discussion of these phenomena in the hyperbolic setting. In [54], a polyhedral realization result for all pairs of cone-metrics satisfying the appropriate curvature condition is derived. Furthermore, the rigidity is established, provided that the cone-metrics are sufficiently small in some sense. In [51], the corresponding hyperbolic realization statement is obtained, and the rigidity is shown in a generic case. Luckily, the mentioned singularities lie outside of the scope of the present article, since they happen away from the Fuchsian locus.
Note that both this paper and [54] exploit the fact that anti-de Sitter geometry is a subgeometry of projective geometry. However, the other tools are quite different. In the present article, we apply the infninitesimal Pogorelov map and obtain local rigidity near the Fuchsian locus. In [54], the technique of geometric transitions is employed, which allows to deduce global rigidity for small enough metrics on the boundary.
@@ 1.4. Our results
There exists a neighborhood $U$ of $\mathscr{M}_{0}$ in $\mathscr{M}$ and an open subset $U^{\prime}\subset\mathscr{S}\times\mathscr{S}$, which contains the diagonal, such that $\mathscr{I}|_{U}$ is a homeomorphism from $U$ to $U^{\prime}$.
@@
For each of those AdS structures in $\mathscr{M}(V)$, the induced metric on $\partial M$ is hyperbolic, with cone-singularities of angle larger than $2\pi$ at the points of $V$. We consider $V_{\pm}$ as subsets of $S$ and denote by $\mathscr{S}(V_{\pm})$ the spaces of hyperbolic metrics with cone-singularities at $V_{\pm}$ respectively of angles $2\pi$. We have the induced metric map $\mathscr{I}_{V}:\mathscr{M}(V)\rightarrow\mathscr{S}(V_{+})\times\mathscr{S}(V_{-})$. We then have the following results.
@@
There exists a neighborhood $U$ of $\mathscr{M}_{0}(V)$ in $\mathscr{M}(V)$ and an open subset $U^{\prime}\subset\mathscr{S}(V_{+})\times\mathscr{S}(V_{-})$, which contains the diagonal, such that $\mathscr{I}_{V}|_{U}$ is a homeomorphism from $U$ to $U^{\prime}$.
@@ 1.5. A word on the proofs
Let $m\in\mathscr{M}$, and $(s_{+},s_{-})=\mathscr{I}(m)$ be the pair of the induced metrics on its boundary. Let $\dot{m}$ be a non-trivial first-order deformation of $m$. Is the corresponding variation $(\dot{s}_{+},\dot{s}_{-})$ of the induced metrics on the boundary necessarily nonzero?
@@ 1.6. Related results
There is a vast body of related results in this area of geometry. Here we survey some of them that we have not mentioned yet, though our exposition does not aim to be exhaustive. For other versions of such a survey, we refer to [30] by Fillastre–Smith and to [59] of Schlenker.
@@
Question 1 is also equivalent to a question of simultaneous isometric embeddings of two metrics on $S$ into GHMC AdS (2+1)-spacetimes. Another interesting problem is to attempt to embed only one metric into a spacetime with prescribed left or right metric in the sense of Mess. This problem in the smooth case was fully resolved in [63] by Tamburelli (realization) and in [19] by Chen–Schlenker (rigidity). In the polyhedral case, the realization was shown in [54] by Prosanov, as well as the rigidity for sufficiently small metrics. In the Minkowski setting, instead of left or right metric it is natural to prescribe the linear part of the holonomy. In the smooth case, this follows from the work [65] by Trapani–Valli, while in the polyhedral case, it is proved in [31] by Fillastre–Prosanov. Actually, we will employ some of these results in our proof. We note that analogous questions in the de Sitter and in the hyperbolic settings (when the left or right metric is replaced by a conformal structure at infinity) are fully open.
Acknowledgments. The first author is very grateful to François Fillastre, Thomas Körber, Andrea Seppi and Graham Smith for useful discussions and to Michael Eichmair for helpful comments and his support. We are very thankful to the anonymous referee for plenty of valuable remarks.
@@
$M$ 3-manifold diffeomorphic to $S\times[-1,1]$
$\mathcal{S}$ the space of Riemannian metrics on $S$ of curvature <1<-1
$\mathcal{M}$
the space of GHC AdS structures on $M$
@@
$\mathcal{R}$ the space of representations $\pi_{1}S\rightarrow G$ up to conjugation
$\mathcal{E}_{\pm}$
equivariant strictly future-/past-convex embeddings of $S$
into $\mathbb{A}d\mathbb{S}^{3}$ up to isometry
$\mathcal{E}$
pairs of embeddings in $\mathcal{E}_{+}\times\mathcal{E}_{-}$
equivariant with respect to the same representation
$\mathcal{GC}_{\pm}(s)$
the spaces of Gauss–Codazzi fields for a metric $s$ on $S$
@@
We add the subscript $F$ for subspaces of Fuchsian objects, i.e., related to the representations of $\pi_{1}S$ into $G_{F}:={\rm PSL}(2,\mathbb{R})$. We add the superscript 0 when anti-de Sitter space is replaced by Minkowski space. We use the font $\mathscr{S}$, $\mathscr{M}$, etc., for the quotients of $\mathcal{S}$, $\mathcal{M}$, etc., by isotopies. We use the hat: $\widehat{\mathcal{R}}$, $\widehat{\mathcal{E}}$, etc., for spaces not up to conjugation/isometries.
@@ 2.1. Anti-de Sitter geometry
The induced metric is a Lorentzian metric of constant curvature 1-1.
We will, however, consider as the main model for anti-de Sitter 3-space the projective quotient of $\mathbb{H}^{2,1}$. We will denote it by $\mathbb{A}d\mathbb{S}^{3}$ and consider it as a subset of $\mathbb{RP}^{3}$. Recall that it is a minimal model space for anti-de Sitter geometry, i.e., any other Lorentzian 3-space of curvature 1-1 with maximal isometry group is a covering of $\mathbb{A}d\mathbb{S}^{3}$, see [17].
We also briefly recall the ${\rm PSL}(2,\mathbb{R})$-model for $\mathbb{A}d\mathbb{S}^{3}$, which is a special feature of dimension 3. Note that on the space of real $\times2$-matrices the determinant is a quadratic form of signature $(2,2)$. Considering the polarization of the minus determinant as a bilinear form, we can identify this space with $\mathbb{R}^{2,2}$. Under this identification, ${\rm SL}(2,\mathbb{R})$ gets identified with $\mathbb{H}^{2,1}$, and ${\rm PSL}(2,\mathbb{R})$ gets identified with $\mathbb{A}d\mathbb{S}^{3}$.
This model admits a straightforward description of the isometry group of $\mathbb{A}d\mathbb{S}^{3}$. We are interested only in its identity component, which we denote by $G$. Since for $A,B,X\in{\rm SL}(2,\mathbb{R})$ we have $\det(X)=\det(AXB^{-1})$, the left and right multiplications by the elements of ${\rm PSL}(2,\mathbb{R})$ are isometries. This produces a map ${\rm PSL}(2,\mathbb{R})\times{\rm PSL}(2,\mathbb{R})\rightarrow G$. A dimension count shows that this is actually an isomorphism
@@
We fix an orientation and a time-orientation on $\mathbb{A}d\mathbb{S}^{3}$. The latter means that we pick a (nowhere zero) timelike vector field $\xi$ on $\mathbb{A}d\mathbb{S}^{3}$ to define future and past. Namely, for every $x\in\mathbb{A}d\mathbb{S}^{3}$ the subset of timelike vectors of $T_{x}\mathbb{A}d\mathbb{S}^{3}$ has two connected components. The vectors from the same component as $\xi(x)$ are called future vectors, the others are called past.
@@ 2.2. Quasi-Fuchsian representations
To proceed further, we need to recall the basics of Teichmüller theory. The Teichmüller space $\mathscr{T}$ is the space of isotopy classes of hyperbolic metrics on $S$. Denote by $\mathbb{H}^{2}$ the hyperbolic plane and by $G_{F}$ the identity component of its isometry group. Every hyperbolic metric on $S$ is a $(G_{F},\mathbb{H}^{2})$-structure in the sense of Thurston. For an introduction to Thurston’s approach to geometric structures (later we will also apply this framework to anti-de Sitter structures), we refer to Thurston’s notes [64] and an accompanying paper [18]. Provided that $S$ and $\mathbb{H}^{2}$ are considered oriented, the holonomies of orientation-preserving hyperbolic structures form a connected component in the representation space of $\pi_{1}S$ to $G_{F}$. We call such representations Fuchsian and denote by $\mathcal{R}_{F}$ the quotient of this component by conjugation of $G_{F}$, the component itself is denoted by $\widehat{\mathcal{R}}_{F}$. In terms of notation, we will not distinguish between representations and their classes. It is well-known that the map sending a hyperbolic structure to its holonomy provides an identification $\mathscr{T}\cong\mathcal{R}_{F}$. See, e.g., the book of Farb–Margalit [24] for an introduction.
Consider $G_{F}\cong{\rm PSL}(2,\mathbb{R})$ as the diagonal subgroup of $G\cong{\rm PSL}(2,\mathbb{R})\times{\rm PSL}(2,\mathbb{R})$. We call a representation $\rho:\pi_{1}S\rightarrow G$ quasi-Fuchsian if it is a deformation of a Fuchsian representation. By definition, quasi-Fuchsian representations form a connected component of a representation space. We denote by $\mathcal{R}$ the quotient by conjugation of $G$, the component itself is denoted by $\widehat{\mathcal{R}}$. The projections of a quasi-Fuchsian representation to both factors in (2.1) are discrete and faithful. Hence, we have $\widehat{\mathcal{R}}\cong\widehat{\mathcal{R}}_{F}\times\widehat{\mathcal{R}}_{F}$ and
@@ 2.3. Globally hyperbolic anti-de Sitter spacetimes
An anti-de Sitter (2+1)-spacetime is a 3-manifold equipped with an anti-de Sitter metric, i.e., with a Lorentzian metric of curvature 1-1. Provided that it is oriented and time-oriented, it is equipped with a $(G,\mathbb{A}d\mathbb{S}^{3})$-structure in the sense of Thurston. Hence, every anti-de Sitter (2+1)-spacetime $M$ is equipped with a holonomy representation $\rho:\pi_{1}M\rightarrow G$ and a $\rho$-equivariant developing map $f:\widetilde{M}\rightarrow G$, which are determined up to an isometry of $\mathbb{A}d\mathbb{S}^{3}$. A (2+1)-spacetime is globally hyperbolic if it admits a Cauchy surface, i.e., a surface that is intersected exactly once by every inextensible causal curve. All Cauchy surfaces in a spacetime are homeomorphic to each other. See, e.g., the book [49] by O’Neill for an introduction to Lorentzian geometry. By a result of Geroch [33], any globally hyperbolic (2+1)-spacetime admits a parametrization $S\times\mathbb{R}$, in which all fibers $S\times\{r\}$ are Cauchy surfaces.
A globally hyperbolic spacetime $N$ is maximal if every isometric embedding $N\rightarrow N^{\prime}$ that sends a Cauchy surface of $N$ to a Cauchy surface of $N^{\prime}$ is onto. It is Cauchy compact if its Cauchy surfaces are compact. Recall that we refer to globally hyperbolic maximal Cauchy compact anti-de Sitter spacetimes as GHMC AdS for short. Any GHMC (2+1)-spacetime $N$ is diffeomorphic to $S\times\mathbb{R}$, and hence $\pi_{1}N=\pi_{1}S$. Let now $N$ be a manifold diffeomorphic to $S\times\mathbb{R}$, and let $\mathscr{N}$ be the space of isotopy classes of GHMC AdS structures on $N$, with the topology induced by uniform $C^{\infty}$-convergence of developing maps on compact subsets of $\widetilde{N}$. Mess showed in [46] that the holonomy of every GHMC AdS (2+1)-spacetime is quasi-Fuchsian. Furthermore, it follows from his work that the map $\mathscr{N}\rightarrow\mathcal{R}$ sending a GHMC AdS structure to its holonomy is a homeomorphism. Any spacelike surface $\Sigma\subset\mathbb{A}d\mathbb{S}^{3}$ invariant with respect to $\rho\in\mathcal{R}$ belongs to the domain of dependence $D(\rho)\subset\mathbb{A}d\mathbb{S}^{3}$, a convex open $\rho$-invariant set, whose quotient by $\rho$ produces the respective GHMC AdS spacetime.